OpenStudy (anonymous):

help please . If y varies directly as x and y = 14 when x = 7, find x when y is 30.

OpenStudy (anonymous):

The answer I came up with is X=15 is this correct

OpenStudy (anonymous):

y1/y2 =x1/x2 or vice versa 30/14=x1/7 therefore you are correct x1=15

OpenStudy (anonymous):

ok thnx have fun now ... :D